Financial Learning Systems has 1.8 million shares of common stock outstanding and 63 comma 939 shares of preferred stock.​ (The preferred pays annual cash dividends of ​$5.89 a​ share, and the common pays annual cash dividends of 32 cents a​ share.) Last​ year, the company generated net profit​ (after taxes) of $ 5 comma 765 comma 438. The​ company's balance sheet shows total assets of ​$73 ​million, total liabilities of $ 33 ​million, and $ 4 million in preferred stock. The​ firm's common stock is currently trading in the market at ​$42.76 a share.

a. Given the preceding​ information, find the​ EPS, P/E​ ratio, and book value per share.

b. What will happen to the price of the stock if EPS rises to ​$4.34 and the​ P/E ratio stays where it​ is? What will happen if EPS drops to ​$1.74 and the​ P/E doesn't​ change?

c. What will happen to the price of the stock if EPS rises to ​$4.34 and the​ P/E jumps to 20.3 times​ earnings?

d. What will happen if both EPS and the​ P/E ratio droplong dash to ​$1.74 and 8.4 times​ earnings, respectively? e. Comment on the effect that EPS and the​ P/E ratio have on the market price of the stock.
Answer #1

common stock = 1,800,000
preferred stock = 63,939, pays dividend of $0.32 per share
PAT = $5,765,438

a) EPS = (PAT - preferred dividend)/common stock =(5,765,438-63,939*0.32)/1,800,000= $3.19
P/E = Price / EPS = 42.76/3.19 = 13.4
To calculate BPS, we need to calculate book value of equity
BV = Assets - Liabilities - Preferred stock = 73-33-4 = 36 Million
BPS = $36 Mn/ 1.8 Mn = $2

b) P/E remains constant at 13.4
if EPS = 4.34, Price = EPS*P/E = 4.34*13.4 =  58.16
if EPS = 1.74, Price = 1.74*13.4 = 23.32

c) Price = EPS*P/E = 4.34*20.3 = 88.1

d) Price = EPS*P/E = 1.74*8.4 = 14.62

e) Higher EPS leads to higher stock price and vice versa. Similarly, higher EPS leads to higher stock price and vice versa